OpenStudy (anonymous):

What is the point-slope equation of the line with slope 1/4 that goes through the point (-2, 8)?

OpenStudy (anonymous):

y = mx+b equation 1: y = (1/4)x+b // find b by plugging in (-2,8) for (x,y). // plug b back into equation 1 and then you have the answer.

OpenStudy (anonymous):

y + 8 = -1/4 (x-2) ?!!

OpenStudy (anonymous):

@tomo

OpenStudy (anonymous):

y = (1/4)x+b 8 = (1/4)(-2) + b 8 = -1/2 + b 8+1/2 = b b = 17/2 just replace b in equation 1 with 17/2